Round off each of the measurements to two significant figures.

Express your answers using two significant figures.

A. 0.00357925 cm

B. 8767 m
1
Expert's answer
2020-12-28T04:40:23-0500

A. 0.00357925 cm

= 3.57925×10^–3 cm

When we take two significant figures then

0.00357925 cm = 3.6×10^–3 cm


B. 8767 m

= 87.67×10^2

When we take two significant figures then

8767 m = 88×10^2 m